If two opposite sides of a square are increased by 15 meters and the other sides are decreased by 13 meters, the area of the rectangle that is formed is 60 square meters. Find the area of the original square.
Answer by jim_thompson5910(35256)

Let x = original side

x goes to x+15 for one pair of sides while the other pair become x-13

So

A = LW

60 = (x+15)(x-13)

(x+15)(x-13) = 60

x^2 + 2x - 195 = 60

x^2 + 2x - 195 - 60 = 0

x^2 + 2x - 255 = 0

(x+17)(x-15) = 0

x+17 = 0 or x-15 = 0

x = -17 or x = 15

Toss out the negative solution to get x = 15 as the only solution.

So the square has a side length of 15 meters.

The area of this square is 15^2 = 15*15 = 225 square meters.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------